the average number of automobiles per minute stopping for a gas at a particular service station along the coastal road is 3. what is a probability that in any given minute more than two automobile will stop for gas?


1
Expert's answer
2021-03-09T16:50:16-0500

We have that

"\mu=3"

x = 2

This follows Poisson distribution

The Poisson probability can be calculated by the formula:

"P(x,\mu)=\frac{e^{-\mu}\mu^x}{x!}"

Need to find "P(x>2,3) = 1 - P(x\le2,3)"

where "P(x\le2,3)= P(0,3)+P(1,3)+P(2,3)"

"P(0,3)=\frac{e^{-3}3^0}{0!}=0.05"

"P(1,3)=\frac{e^{-3}3^1}{1!}=0.15"

"P(2,3)=\frac{e^{-3}3^2}{2!}=0.22"

"P(x>2,3) =1-0.05-0.15-0.22=0.58"


Answer: 0.58